OpenStudy (anonymous):

What is e^-ln(x+1)? Is it -x-1 or 1/(x+1)

OpenStudy (anonymous):

hint \[-\ln(a)=\ln(\frac{1}{a})\]

OpenStudy (anonymous):

thank you